### v2

# Non-Linearities in International Spillovers of the ECB’s Monetary Policy

Abstract
We investigate the presence of sign and size non-linearities in the impact of the European Central Bank’s Anti-Fragmentation Policy on non-ERM II, EU countries. After identifying three orthogonal monetary policy shock using the method of Fanelli and Marsi [2022], we then select an optimal specification and estimate both linear and non linear impulse response functions using local projections (Dufour and Renault [1998], Gonçalves et al. [2021]). The choice of non-linear transformations to separate sign and size effects is based on Caravello and Martinez-Bruera [Working Paper, 2024]. Lastly we compare the linear model to the non-linear ones using a battery of Wald tests and find significant evidence of sign non-linearities in the international spillovers of ECB policy.
JEL: F10, F33, F40, G32.
1 Introduction
In a world where financial markets are connected internationally, monetary policy spills over to neighbouring countries. In the context of American monetary policy the question, has been studied originally by Fleming [1962], Mundell [1963] and Dornbusch [1976]. More recent papers such as Jarociński [2022] and Camara [R&R, 2023] use a high-frequency identification strategy to look at the spillovers to the European Union and Emerging Markets respectively. The European central bank has also been shown to produce international spillovers. For example, Jarociński [2022] finds that ECB conventional monetary policy produce small spillovers to the US and Miranda-Agrippino and Nenova [2022] finds spillovers from unconventional monetary policy.
In this paper we study the impact of identified monetary policy shocks from the European Central Bank on EU countries that are not part of the European Exchange Rate Mechanism from 2002 to 2019. That is, countries in the EU but outside the Euro-zone, that have a free floating currency. This is related to Hajek and Horvath [2018], which studies the spillovers to non euro-zone eu countries. However, instead of proxying conventional and unconventional monetary policy with the shadow rate, we use the identification strategy of Fanelli and Marsi [2022] to recover three orthogonal monetary policy shocks. 1) A pure monetary policy shock, which represents the actual effect of monetary policy (borrowing costs). 2) An information shock, which represents the information about the current and future state of the economy that can be inferred from monetary policy (overheating, stagnation). 3) A spread shock, which captures anti-fragmentation policies of the ECB. We are interested in the presence of non-linearities in the effects of these monetary policy shocks motivated by the fact the ECB’s main refinancing rate was close to the zero lower bound for nearly a decade (2012-2022).
In section 1.1, we go over the European Central Bank’s anti-fragmentation policy and the implication of the zero-lower bound on the non-linear behavior of monetary policy. In section 2, we give a detailed description of our data. Section 3, goes over our econometric methodology, including the postulated underlying data generating process, the linear and non-linear models and the model selection. Section 4, provides the results of the impulse response functions estimation and lastly, section 5, provides inference on the non-linear specifications, both about the presence of non-linearities and the significance of the IRF itself .
1.1 Anti-Fragmentation and the Zero Lower Bound
Fragmentation in the sovereign debt markets in the EU refers to the phenomenon where bond spreads between countries with higher levels of debt to GDP (Greece, Portugal, Spain, Italy) and Germany increase in periods of monetary contractions (or transition from expansion to a more neutral stance). For example, the Portuguese-German spread in the 10-year bond rate reached nearly $14\%$ and more famously, the Greek-German spread reached $32\%$ [^2] https://www.europarl.europa.eu/RegData/etudes/STUD/2022/733984/IPOL_STU(2022)733984_EN.pdf. Sovereign spreads indicate different market expectations of default risks between countries. Therefore, their presence suggests that the European financial markets are not fully integrated. Contractionary monetary policy from the ECB will affect riskier countries more adversely. Meaning, the transmission of monetary policy is harmed (Wyplosz [2022]). To combat this, the ECB began intervening in sovereign debt markets, starting in 2010, with the securities market program (SMP), as detailed in Dabrowski [2022]. With this program, Euro-zone countries’ central banks could purchase government assets to lower the cost of borrowing in their respective sovereign debt markets. In Greece, the program led to a decrease in the 5 year rate of up to $0.21\%$ for every billion euros in purchases without however, increasing the quantitaty of money in circulation as: “Any additional monetary base created […] was to be sterilised by the ECB in its open-market operations” (Dabrowski [2022]). Other programs eventually replaced the SMP such as the Outright Monetary Transaction (OMT) program and the Emergency Liquidity assistance (ELA), all with the same goal of reducing fragmentation risk. In July 2022, the ECB unveiled the Transmission Protection Instrument (TPI), a program allowing the purchase of governmental securities with maturities of one to ten years [^3] https://www.ecb.europa.eu/press/pr/date/2022/html/ecb.pr220721~973e6e7273.en.html.
Estimating the spillovers of anti-fragmentation and conventional monetary policy is complicated by the fact that since 2012, the ECB’s refinancing rate is near the zero-lower bound (ZLB). Recent research suggests that dynamic stochastic general equilibrium models should be estimated with non-linear methods to replicate outcomes near the ZLB Atkinson et al. [2020], Fernández-Villaverde et al. [2015], Wu and Xia [2016]. Other papers, such as Caggiano et al. [2017], use non-linear transformation of variables in vector auto regressive models to study the behavior of the economy near the ZLB. Our paper is therefore related in the sense that we use the ZLB to justify investigating the presence of non-linearities in the effects of monetary policy.
2 Data
2.1 Country Specific Variables
We build an unbalanced panel of European Union countries at the monthly frequency from the period of January 2002 to February 2023 (Appendix D, presents the results for the pre-pandemic sub-sample January 2002 to December 2019). The panel includes countries that satisfies the following conditions: 1) the country is in the European Union, 2) the country has a free floating currency that is not the Euro, that is, countries that are not part of the European Exchange Rate Mechanism (ERM 2). There are 5 countries that satisfy these conditions: Romania, Poland, Hungary, Czechia and Sweden. Table 1, shows the composition of the unbalanced panel.
Country Entry Date Exit Date
Czechia 2002M01 2023M02
Hungary 2002M01 2023M02
Poland 2002M01 2023M02
Romania 2005M04 2017M09
Sweden 2002M01 2017M05
Table 1: Composition of the Panel
For each country we first collect 4 endogenous variables from the IMF dataset: a seasonally adjusted industrial production index, the consumer price index, the unemployment rate and a measure of the long term interest rate. We remove the seasonality of unemployment by projecting it on the orthogonal complement of the column space of a matrix whose columns are monthly dummies.
We then add a trade-valued real exchange rate to every country in the panel, which is sourced from Bruegel [^4] https://www.bruegel.org/publications/datasets/real-effective-exchange-rates-for-178-countries-a-new-database . This measure is a weighted average based on a basket of 120 trading partners of each country.
2.2 Euro Area Control Variables
We collect 4 control variables for the euro area. An aggregate of the consumer price index is taken from the IMF dataset. Seasonally adjusted unemployement rate and industrial production is taken from the OECD dataset. Lastly, we also include the 120 countries trade valued real exchange rate for the Euro from Bruegel. However, the trading partner weights are computed based on the first 12 countries of the Eurozone. Our panel’s countries are not part of this list.
2.3 Identified Shocks
To extract our identified shocks we follow the three-shocks methodology of Fanelli and Marsi [2022]. They extract three exogenous shocks from high-frequency movements of prices on a portfolio of European financial assets. These shocks are 1) a pure monetary policy shock, 2) a central bank information shock and 3) a spread shock attributed to the ECB’s anti-fragmentation policy. The high-frequency surprises are calculated as the difference between the median quote price in a 10 minute window 20 minute before the start of an ECB press conference and the median quote in a 10 minute window 55 minutes after the start of the conference. This difference will be interpreted as a “surprise” unanticipated by the market. They postulate that the three exogenous shocks are related to these high-frequency surprises by the static factor model:
$\Delta_{t}=$ $\Lambda f_{t}+e_{t}$
[…]
$\Delta OIS_{iy,t}$, is the surprise in the price of overnight indexed swaps on the “i” years interest rate, $\Delta IT_{iy,t}$ is the surprise on the rate of Italian bonds of “i” year maturity and $\Delta STOXX50$ is the surprise in the STOXX50 a Eurozone stock exchange index. These variables come from the EA-MPD dataset established by Altavilla et al. [2019].
The matrix of factor loadings $\Lambda$ and the factors $f_{t}$ are jointly estimated by maximum likelyhood. However, if $\Lambda^{*}$ and $f_{t}^{*}$ are the maximum likelyhood estimates, post multiplying $\Lambda$ by an orthonormal matrix $Q$ and premultiplying $f_{t}$ by the transpose of Q also satisfies the same equation. Therefore, the authors introduce sign restrictions on the factor loadings which allow set identification of theoretically sound factors. $+$ and -- indicates positive and negative co movement respectively. $\bullet$ represents the absence of restrictions and --- - indicates that the magnitude of the movement must be greater than all other factors. To sum up, the pure monetary policy shock ($f_{mp}$) coincides with higher rates, decreases in the stock index and an increase in the spreads (due to reasons outlined in section 1.1). A positive information shock ($f_{inf}$) announces to the markets that the economy is overheating. Therefore, it is associated with an increase in the stock indexes and a decrease in the spreads. Lastly, a positive spread shock ($f_{spread}$) should lower the spreads and increase the stock indexes.
We estimate the factor model using data from january 2002 up to November 2023, the latest date available in the EA-MPD dataset at the time. As in Fanelli and Marsi [2022], we randomly generate orthonormal matrices Q. If $\Lambda Q$ satisfies the sign restrictions, it is kept. Denote $\Lambda^{*}Q^{*}$ to be the median rotated factor loading to satisfy the sign restrictions. $(Q^{*})^{\prime}f_{t}$ are our factors.
These three factors are the identified exogenous shock whose impact we will study. We transform them to monthly time series by assigning to each month the value of the shock from a conference that happened during that month. Months without any ECB conferences have values of 0. Exceptionally august 2006 had two ECB conferences while September 2006 had none. Since, the second conference was on the 31st of august we assign it to the month of september 2006.
In the next section we will outline our econometric methodology.
3 Econometric methodology
In this section we outline the main components of our econometric methodology. We start with a description of the hypothesized underlying structural model. We then describe the estimation method, the model selection and lastly, our strategy to recover possible sign and size effects of the identified shocks.
3.1 The Structural Model
We postulate that the data generating process, for each country j, is a structural vector auto-regressive model of the following form (similar to the one studied in Gonçalves et al. [2021] and the VARX of Dees et al. [2007]):
$A_{0}w_{j,t}=a+A(L)w_{j,t-1}+C(L)f(x_{t})+\eta_{j,t}$ (1)
Where $w_{j,t}=\langle x_{t},y_{j,t},z_{t}\rangle$, $x_{t}$ is a vector of identified shocks, $y_{j,t}$ is a vector of country specific endogenous variables and $z_{t}$ are the euro-area endogenous variables.
$A_{0}=\begin{bmatrix}\mathbf{I}&\mathbf{0}&\mathbf{0}\\ -A_{0,21}&\mathbf{I}&-A_{0,23}\\ -A_{0,31}&-A_{0,32}&\mathbf{I}\end{bmatrix}$ (2)
The elements of $y_{j,t}$ and $z_{t}$ do not impact $x_{t}$ contemporaneously but $x_{t}$ does affects $y_{j,t}$ and $z_{t}$ contemporaneously. We do not restrict the contemporaneous relationship between $y_{j,t}$ and $z_{t}$ as we only require a block-recursive DGP to identify the impulse response to a shock in $x_{t}$.
$A(L)=\begin{bmatrix}\mathbf{0}&\mathbf{0}&\mathbf{0}\\ A_{21}(L)&A_{22}(L)&A_{23}(L)\\ A_{31}(L)&\mathbf{0}&A_{33}(L)\\ \end{bmatrix}$ (3)
Where $A_{21}(L)$, $A_{22}(L)$ and $A_{23}(L)$ are polynomials of the lag operator of order $p$, $q$ and $r$ respectively. The first row of sub-matrices of $A(L)$ are zero matrices as our shocks are exogenous and observed.
As in Gonçalves et al. [2021], we assume the following structure for $C(L)$:
$C(L)=\begin{bmatrix}\mathbf{0}\\ C_{2}(L)\\ C_{3}(L)\end{bmatrix}$ (4)
By their construction, the identified shocks are uncorrelated between themselves and have unit variances.
$\eta_{j,t}=\langle u_{t},\epsilon_{j,t},v_{t}\rangle\sim i.i.d\left(\begin{bmatrix}0\\ 0\\ 0\end{bmatrix},\begin{bmatrix}\mathbf{I}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\Sigma_{2,2}&\Sigma_{2,3}\\ \mathbf{0}&\Sigma_{3,2}&\Sigma_{3,3}\\ \end{bmatrix}\right)$ (5)
The object of interest in our analysis is the impulse response to a shock in $u_{t}$ to $y_{t}$. Since, the shocks are observed, estimating the impulse response conditional on the sample is akin to estimating the unconditional impulse response function (Gonçalves et al. [2021]).
3.2 The Linear Local Projection and Model Selection
We estimate the impulse responses of interest using the method of local projections (Dufour and Renault [1998] and Jordà [2005]). In the first baseline specification we treat $C_{2}(L)$ as a zero matrix, completely ignoring non-linearities. We refer to this as the linear specification which yields the following local projection.
$y_{j,k,t+h}=$ $\alpha_{j,k,h}+\sum_{i=0}^{p}\psi_{j,h,i}x_{t-i}+\sum_{i=1}^{q}\phi_{j,h,i}y_{k,t-i}+\sum_{i=1}^{r}\beta_{j,h,i}z_{t-i}+I_{1}\beta_{j,h}t+I_{1}I_{2}\gamma_{j,h}t^{2}+\epsilon_{j,k,t},$ (6)
$h\in\{0,...,H\}$ $\quad k\in\{1,...,5\}\quad j\in\{1,...,5\}\quad\langle p,q,r\rangle\in\mathbf{N}^{3}\quad\langle I_{1},I_{2}\rangle\in\{0,1\}^{2}\quad t\in\{0,...,T\}$
The unconditional impulse response function for a shock $\delta$ will then be:
$IRF_{j,h,\delta}=\psi_{j,h,0}\delta$
Notation, in equation (6), j is the subscript for the country specific endogenous variable, k the country, t the time, h the horizon and i the lag. That is,$y_{j,k,t+h}$ corresponds to the endogenous variable j for country k at time t+k, while $y_{k,t-i}$ is the 5 by 1 vector of endogenous variable for country k at time t-i. $\alpha_{j,k,h}$ is a equation-country-horizon specific intercept. $x_{t-i}$ is the 3 by 1 vector of shocks and $z_{t-i}$ is the 4 by 1 vector of Euro-Area endogenous variables. $x_{t-i}$ and $z_{t-i}$ are common to every country and every equation j. $t$ and $t^{2}$ are the linear and quadratic time trends and $I_{1}$ and $I_{2}$ are either 0 or 1 and will be used in model selection.
Model Selection We select the optimal number of lags p, q and r as well as the presence of a linear and quadratic time trend at every horizon $h$ using the information criteria of Akaike [1973].
Denote by $\hat{u}^{h}(p,q,r,I_{1},I_{2})$ the residuals from estimating equation (6) with OLS. Define the sets: $A=\{2,3,4,5,6\}$ and $B=\{0,1\}$. We select $p^{*}(h)$, $q^{*}(h)$, $r^{*}(h)$, $I_{1}^{*}(h)$ and $I_{2}^{*}(h)$ to solve the following problem at each $h\in\{0,..,H\}$:
$\underset{\{p,q,r\}\in A^{3},\{I_{1},I_{2}\}\in B^{2}}{\text{max}}n*ln((\hat{u}^{h}(p,q,r,I_{1},I_{2})^{\prime}\hat{u}^{h}(p,q,r,I_{1},I_{2}))/n)-2(p+q+r+I_{1}+I_{1}I_{2}+5)$ (7)
Appendix A, provides the detailed results of this model selection, while appendix B, provides the results if we instead used the BIC criterion of Schwarz [1978] .
We then estimate both the standard linear local projection and the local projection augmented with non-linear transformation of the shocks using the optimal selection $p^{*}(h)$, $q^{*}(h)$, $r^{*}(h)$, $I_{1}^{*}(h)$ and $I_{2}^{*}(h)$.
3.3 Uncovering Non-Linearities
To investigate the presence of non-linearities in the impulse response functions we augment our local projections with non-linear transformation of the shocks. That is, we specify a $f(x)$ and we let $C_{2}(L)$, be non-zero. We are interested in two types of non-linearities. Sign based non-linearity and Size based non-linearity. Our choice of function $f(x)$ to disentangle the two effects will be taken from Caravello and Martinez-Bruera [Working Paper, 2024]. We then compute the impulse responses using the plug in non linear local projection method of Gonçalves et al. [2021].
Caravello and Martinez-Bruera [Working Paper, 2024] define Sign Non-Linearities as when the magnitude of the effect of a shock depending on the sign of the shock. To estimate the sign effect they propose augmenting the local projection with the absolute value of the shock: $f(x_{t})=|x_{t}|$. They demonstrate that this function will annihilate size non-linearities and recover only the sign effect. For example, a positive shock will have an immediate impact of: $\Gamma_{j,0,0}+\psi_{j,0,0}$, while for a negative shock it will be: $\Gamma_{j,0,0}-\psi_{j,0,0}$. This result depends on the symmetry of the distribution of shocks around 0, which is discussed in appendix C. We call the following the sign specification local projection:
$y_{j,k,t+h}=$ $\alpha_{j,k,h}+\sum_{i=0}^{p^{*}(h)}\psi_{j,h,i}x_{t-i}+\sum_{i=0}^{p^{*}(h)}\Gamma_{j,h,i}|x_{t-i}|+\sum_{i=1}^{q^{*}(h)}\phi_{j,h,i}y_{k,t-i}$
$+\sum_{i=1}^{r^{*}(h)}\beta_{j,h,i}z_{t-i}+I_{1}^{*}\beta_{j,h}t+I_{1}^{*}(h)I_{2}^{*}(h)\gamma_{j,h}t^{2}+\epsilon_{j,k,t}$ (8)
$h\in\{0,...,H\}$ $\quad k\in\{1,...,5\}\quad j\in\{1,...,5\}\quad t\in\{0,...,T\}$
Where $p^{*}(h),q^{*}(h),r^{*}(h),I_{1}^{*}(h),I_{2}^{*}(h)$ are the optimal parameters selected in for the linear specification.
Size Non-Linearities occur when the marginal effect of the shock depends on the magnitude of the shock. For example, if increasing the interest rate twice by $1\%$ does not have the same average effect as increasing it once by $2\%$. Size non-linearities are trickier to report as the scaled impulse response depends on the magnitude of the shock. The authors recommend picking a threshold $\bar{b}$ on the shock’s magnitude to partition them into small and large. The non-linear transformation of the shocks will then be: $f(x_{t})=\mathbf{I}\{x_{t}\leq-\bar{b}\}(x_{t}+\bar{b})+\mathbf{I}\{x_{t}\geq\bar{b}\}(x_{t}-\bar{b})$. This non-linear transformation will capture the average marginal effect for the part of the shock above the magnitude threshold. We denote this specification the size local projection:
$y_{j,k,t+h}=$ $\alpha_{j,k,h}+\sum_{i=0}^{p^{*}(h)}\psi_{j,h,i}x_{t-i}+\sum_{i=0}^{p^{*}(h)}\Gamma_{j,h,i}\left[\mathbf{I}\{x_{t-i}\leq-\bar{b}\}(x_{t-i}+\bar{b})+\mathbf{I}\{x_{t-i}\geq\bar{b}\}(x_{t-i}-\bar{b})\right]$
$+\sum_{i=1}^{q^{*}(h)}\phi_{j,h,i}y_{k,t-i}+\sum_{i=1}^{r^{*}(h)}\beta_{j,h,i}z_{t-i}+I_{1}^{*}\beta_{j,h}t+I_{1}^{*}(h)I_{2}^{*}(h)\gamma_{j,h}t^{2}+\epsilon_{j,k,t}$ (9)
$h\in\{0,...,H\}$ $\quad k\in\{1,...,5\}\quad j\in\{1,...,5\}\quad t\in\{0,...,T\}$
[…]
As the impulse response function is non-linear, its shape depends on the magnitude of the shock, $\delta$, and the initial value of $x_{t}$. For example, with a size non-linearity if $x_{t}=0.5$ and $\delta=1$ the response will be different than if $x_{t}=2$ and $\delta=1$. Gonçalves et al. [2021] provides us with a method to report the unconditional impulse response using a plug in local projection estimator. We outline their procedure here:
1) We estimate the local projections in the standard way which yields for each equation j and each horizon h a pair $\hat{\psi}_{j,h,0}$ and $\hat{\Gamma}_{j,h,0}$.
2) We estimate the quantity: $A_{\delta}=E(f(x_{t}+\delta)-f(x_{t}))$ with $\hat{A}_{\delta}=\frac{1}{T}\sum_{t=1}^{T}\left[f(x_{t}+\delta)-f(x_{t})\right]$ for each of the three shocks.
3) The impulse response function will be: $IRF(h,j)=\hat{\psi}_{j,h,0}\delta+\hat{\Gamma}_{j,h,0}\hat{A}_{\delta}$.
This method relies on the assumption that the shocks are not serially correlated. Estimating AR(1) models of each shocks yields non-significant coefficients that are less than 0.1.
Lastly we also report conditional impulse responses for the sign non-linearity specification. These are the impulse responses conditional on $x_{t}$ being non-negative (non-positive) and being disturbed by $\delta$ that is positive (negative). We compute them as: $IRF(h,j|x_{t}\geq 0,\delta>0)=\hat{\Gamma}_{j,h,0}+\hat{\psi}_{j,h,0}$ and $IRF(h|x_{t}\leq 0,\delta<0)=\hat{\Gamma}_{j,h,0}-\hat{\psi}_{j,h,0}$ respectively.
The inference on the non-linear models will be two-pronged. We first look at the significance of the coefficient on the non-linear transformation of the shock, to establish the presence of non-linearities. We then test the significance of both the unconditional and conditional IRFs, to determine if they are non-zero. The inference on the non-linear models will be discussed in great details in section 5.
4 IRF Estimation Results
In this section we provide the estimated impulse response functions. We start with the linear specification and move to both the sign and size non-linear specifications.
4.1 The Linear Model
Figure 1: Results from the linear specification (6)
[…]
The horizontal axis represents the months since the shock (h). The rows are in order: the real exchange rate, the unemployment rate, the consumer price index, the index of industrial production and the long term interest rate. The columns are in order: the pure monetary policy shock, the information shock and the spread shock. The real exchange rate, consumer price index and industrial production index are in log points, while the unemployment rate and long term interest rate are in percentages. The dark blue lines represent the 1 standard deviation confidence interval while the light blue are $90\%$. The standard errors are computed using the method of Arellano [1987] with the “Jackknife” $HC_{3}$, HCCME..
Figure 1, provides the estimated impulse response functions of the linear specification. The pure monetary policy shock has significant, but temporary, effects on the real exchange rate and the CPI. The information shock has an quick but temporary impact on the real exchange rate and a delayed (and less significant) impact on CPI. Lastly, the spread shock has a very significant negative impact on the unemployment rate and positive impact on industrial production. For nominal variables, impact on CPI, the REER and the long term rate is positive and significant.
4.2 The Non-Linear Model, Sign Non-Linearities
Figure 2: Impulse responses of the sign specification à la Gonçalves et al. [2021]
[…]
The red line is the original linear impulse response function, the blue line is the Gonçalves et al. [2021] non-linear IRF. The horizontal axis represents the months since the shock (h). The rows are in order: the real exchange rate, the unemployment rate, the consumer price index, the index of industrial production and the long term interest rate. The columns are in order: the pure monetary policy shock, the information shock and the spread shock. The real exchange rate, consumer price index and industrial production index are in log points, while the unemployment rate and long term interest rate are in percentages.
In figure 2, we plot in blue our estimate of the unconditional impulse response function to a 1 standard deviation disturbance in the sign-effect non-linear specification. In red is the original linear impulse response function to a 1 standard deviation shock. When comparing both we see that for most graphs the two IRFs nearly coincide, implying for them that on average both specification will estimate very similar IRFs. However, the monetary shock on the long term rate, is a lot stronger (temporarily) starting from horizon 5, when estimated with the sign-specification. For the information shock, the linear model estimates an effect on unemployment and industrial production that is close to zero, while the non-linear model estimates a much larger and positive IRF for unemployment and a negative IRF for industrial production. This implies that the effect of the information shock on real variables contains a sign effect. Lastly, the non-linear model estimates a larger, but temporary effect of the spread shock on industrial production, the real exchange rate,the consumer price index and the long term rate.
Figure 3: Conditional IRFs of the sign-effect specification
[…]
The rows are in order: the real exchange rate, the unemployment rate, the consumer price index, the index of industrial production and the long term interest rate. The columns are in order: the pure monetary policy shock, the information shock and the spread shock. The real exchange rate, consumer price index and industrial production index are in log points, while the unemployment rate and long term interest rate are in percentages.
Figure 3, compares two conditional impulse response functions estimated with the sign-effect specification. These are:
[…]
The blue IRFs are very similar to the unconditional Gonçalves et al. [2021] IRFs as the standard deviation and the estimated $\hat{A}$ are both close to 1. We can see that for monetary policy shocks, a (conditional) positive shock has a large negative effect on REER and a large positive effect on the long term rate, while a negative shock has a more restrained effect on both. For industrial production, it is the opposite, negative monetary policy shocks have more negative impacts. Positive information shocks have stronger permanent effects on unemployment. For the spread shock the main difference is in its effect on the unemployment rate. Positive and negative shocks have opposite marginal effects. A positive spread shocks also has a much larger effect on the long term rate. All other conditional Impulse response functions are similar when comparing negative and positive shocks.
4.3 The Non-Linear Model, Size Non-Linearities
Figure 4: Impulse responses of the size specification à la Gonçalves et al. [2021]
[…]
The rows are in order: the real exchange rate, the unemployment rate, the consumer price index, the index of industrial production and the long term interest rate. The columns are in order: the pure monetary policy shock, the information shock and the spread shock. The real exchange rate, consumer price index and industrial production index are in log points, while the unemployment rate and long term interest rate are in percentages.
Figure 4 plots the Gonçalves et al. [2021] unconditional impulse response functions for the size-effect specification in blue. We set the threshold $\bar{b}$ for each shock to be such that $P(|x_{t}|\leq\bar{b})\approx 0.6$. The shock used in computing of $\hat{A}$ is still 1 standard deviation. In red is the original linear impulse response function. The monetary policy shock has a slightly larger temporary negative effect on the REER for horizon 13 to 22. The information shock has a smaller effect on the unemployment rate, CPI and Industrial production when it is estimated with the non-linear specification. The spread shock has a smaller effect on CPI and a larger effect on the long term rate at longer horizons.For all other impulse response functions, the linear and non-linear specification coincide. This suggests that there are no important size effects for a 1 standard deviation shock.
Figure 5: Impulse responses of the size specification à la Gonçalves et al. [2021]
[…]
The magenta, orange, red, green and blue line are for shocks that are 0.5, 0.75, 1, 1.25 and 1.5 times the standard deviation respectively. The rows are in order: the real exchange rate, the unemployment rate, the consumer price index, the index of industrial production and the long term interest rate. The columns are in order: the pure monetary policy shock, the information shock and the spread shock. The real exchange rate, consumer price index and industrial production index are in log points, while the unemployment rate and long term interest rate are in percentages.
Figure 5 presents an alternative way of illustrating the size effect using the method of Gonçalves et al. [2021]. Each line is an impulse response of the following form:
$IRF(h,a)=\frac{1}{a}\left(\hat{\psi}_{j,h,0}*a*\sigma+\sum_{t=1}^{T}\frac{1}{T}\left[f(x_{t}+a*\sigma)-f(x_{t})\right]*a*\hat{\Gamma}_{j,h,0}\right)$
Where $a\in\{0.5,0.75,1,1.25,1.5\}$. We can see that for many impulse responses the shape depends greatly on the size of the shock. This suggest that even if we did not see any differences in the impulse response between the linear and non-linear specification for the average shock (1 standard deviation), it could appear for different size shocks. For example, monetary policy shocks on the REER have bigger marginal effects for larger shocks for the first year ($h<12$), but smaller effects for the second year. The information shock’s marginal effect on REER and Unemployment is larger for smaller shocks. Lastly, the spread shock’s marginal effect on the REER, the CPI and Industrial Production is larger for smaller magnitude shocks. Therefore, if there are size effects, they are concave in the magnitude of the shock. It is important to remember that the significance level of the coefficient of the non-linear transformation (and therefore the presence of non-linearities) does not depend on the size of the shock.
5 Inference on the Non-Linear Models
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24 25
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 2: Significance of the Non-Linear Component of the Sign Effect Specification
[…]
We report the significance level of the estimated coefficient $\hat{\Gamma}_{j,h,0}$. The non-linear transformation is the absolute value of the shocks. White cells refer to p-values greater than 0.1, yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 2 reports the results of Wald tests (Wald [1943]) of the significance of the coefficient on the non-linear transformation of the shocks. As in Caravello and Martinez-Bruera [Working Paper, 2024], the results of these tests allow us to establish the presence of non-linearities at horizon h. We find multi-horizon significant sign-effects of every type of shock on every outcome variable with the exception of the monetary shock on the long term rate, the information shock on CPI and the long term rate and the spread shock on unemployement. We describe the structure of the test bellow as we will present results on different version of it. We test the hypotheses,
$H_{0}^{1,h}:\hat{\Gamma}^{monetary}_{j,h,0}=0\quad H_{0}^{2,h}:\hat{\Gamma}^{information}_{j,h,0}=0\quad H_{0}^{3,h}:\hat{\Gamma}^{spread}_{j,h,0}=0$
To do so we conduct the Wald test:
$\hat{\beta}_{j,h}=$ $\langle\hat{\psi}^{monetary}_{j,h,0},\hat{\psi}^{information}_{j,h,0},\hat{\psi}^{spread}_{j,h,0},\hat{\Gamma}^{monetary}_{j,h,0},\hat{\Gamma}^{information}_{j,h,0},\hat{\Gamma}^{spread}_{j,h,0},...\rangle$
$R_{1}=$ $\langle 0,0,0,1,0,0,...\rangle,\quad R_{2}=\langle 0,0,0,0,1,0,...\rangle,\quad R_{3}=\langle 0,0,0,0,0,1,...\rangle$
$W_{i,j,h}=$ $(R_{i}\hat{\beta_{j,h}}^{\prime})^{\prime}(R_{i}\hat{\Omega}R_{i}^{\prime})^{-1}(R\hat{\beta_{j,h}}^{\prime})$
[…]
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24 25
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 3: Significance of the Non-Linear Component of the Size Effect Specification
[…]
We report the significance level of the estimated coefficient $\hat{\Gamma}_{j,h,0}$. The non-linear transformation is $f(x_{t})=\mathbf{I}\{x_{t}\leq-\bar{b}\}(x_{t}+\bar{b})+\mathbf{I}\{x_{t}\geq\bar{b}\}(x_{t}-\bar{b})$ with $\bar{b}$ chosen such that $P(|x_{t}|<\bar{b})=0.6$. White cells refer to p-values greater than 0.1, yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 3 reports the result of the same Wald test as in table 2, but on the size effect specification. That is:
$R_{1}=$ $\langle 0,0,0,1,0,0,...\rangle,\quad R_{2}=\langle 0,0,0,0,1,0,...\rangle,\quad R_{3}=\langle 0,0,0,0,0,1,...\rangle$
The results are more muted then with the sign effect. For example, there are very little size effects of any shock on the REER and the unemployement rate for the first 12 months. The information shock has no size effect on industrial production. The persistent and significant non-linearities are that of the monetary shock on the long term rate and the spread shock on the CPI.
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24 25
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 4: Sign-Effect Impulse Response Function à la Gonçalves et al. [2021]
[…]
We report the significance level of the Gonçalves et al. [2021] IRF with the sign effect specification. White cells refer to p-values greater than 0.1, yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 4, reports our attempt at testing the significance of the Gonçalves et al. [2021] with a Wald test. This corresponds to the the following three hypotheses at each horizon,
$H_{0}^{i,h}:\hat{\psi}_{j,h,0}^{i}\delta^{i}+\hat{A}^{i}\hat{\Gamma}_{j,h,0}^{i}=0,\quad i\in\{monetary,information,spread\}$
To do so, we construct the following matrix of linear restrictions:
$R_{1}=$ $\langle\delta^{monetary},0,0,\hat{A}^{monetary}(\delta^{monetary}),0,0,...\rangle$
$R_{2}=$ $\langle 0,\delta^{information},0,0,\hat{A}^{information}(\delta^{information}),0,...\rangle$
$R_{3}=$ $\langle 0,0,\delta^{spread},0,0,\hat{A}^{spread}(\delta^{spread}),...\rangle$
Where: $\delta^{i}$ is the sample standard deviation of shock $i$ and $\hat{A}^{i}=\frac{1}{T}\sum_{t=1}^{T}\left[f(x_{t}^{i}+\delta^{i})-f(x_{t}^{i})\right]$, and recall that for the sign effect specification $f(x)=|x|$. The Wald Statistic is then constructed in the same way as before. This procedure is not exact as it does not take into account the estimation uncertainty in $\hat{A}^{i}$. Although we previously established the presence (or not) of non-linearities, the goal of this exercise is to see if these non-linearities do not “cancel out” the linear effect such that the weighted sum of both is not significant. The Gonçalves et al. [2021] IRF with the sign effect specification is significant for monetary shocks to the REER and the long term rate and the spread shock to CPI, unemployment and industrial production. The sign and magnitude of the impulse response is reported in figure 2.
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24 25
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 5: Size-Effect Impulse Response Function à la Gonçalves et al. [2021]
[…]
We report the significance level of the Gonçalves et al. [2021] IRF with the size effect specification. White cells refer to p-values greater than 0.1, yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 5 reports results of the same test as in table 4, albeit on the size-effect specification, with $\bar{b}$ such that $P(|x_{t}|<\bar{b})=0.6$. That is we test the hypothesis:
$H_{0}^{i,h}:\hat{\psi}_{j,h,0}^{i}\delta^{i}+\hat{A}^{i}\hat{\Gamma}_{j,h,0}^{i}=0,\quad i\in\{monetary,information,spread\}$
The wald statistic is constructed in the same way as in table 4. As when we tested the significance of $\hat{\Gamma}_{j,h,0}^{i}$ alone, we find more muted results. The impulse response function is only significant (for multiple horizons) for monetary shocks to REER and for spread shocks to unemployment.
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24 25
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 6: Conditional IRFs Positive Sign-Effect
[…]
We report the significance level of the conditional IRFs for a positive shock with the sign effect specification. White cells refer to p-values greater than 0.1, yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 6, provides the results for significance tests on the sign-effect, positive-shock, conditional impulse response functions. That is we test the hypotheses:
$H_{0}^{i,h}:\hat{\psi}_{j,h,0}^{i}+\hat{\Gamma}_{j,h,0}^{i}=0,\quad i\in\{monetary,information,spread\}$
To do so we construct the restriction matrix:
$R_{1}=\langle 1,0,0,1,0,0,...\rangle\quad R_{2}=\langle 0,1,0,0,1,0,...\rangle\quad R_{3}=\langle 0,0,1,0,0,1,...\rangle$
We then construct the Wald statistic in the same way. We find that conditional on $x_{t}\geq 0$ and $\delta>0$, there are significant effects of the monetary policy shock to REER, CPI and the Long term Rate, significant effects of the information shock to unemployment at horizon 11 to 18 and significant effects of the spread shocks to CPI, Industry and Unemployment.
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24 25
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 7: Conditional IRFs Negative Sign-Effect
[…]
We report the significance level of the conditional IRFs for a negative shock with the sign effect specification. Yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 7, provides the results for significance tests on the sign-effect, negative-shock, conditional impulse response functions. That is we test the hypotheses:
$H_{0}^{i,h}:\hat{\Gamma}_{j,h,0}^{i}-\hat{\psi}_{j,h,0}^{i}=0,\quad i\in\{monetary,information,spread\}$
Which imply the restriction matrices:
$R_{1}=\langle-1,0,0,1,0,0,...\rangle\quad R_{2}=\langle 0,-1,0,0,1,0,...\rangle\quad R_{3}=\langle 0,0,-1,0,0,1,...\rangle$
From the wald test we find that if $\delta<0$ and $x_{t}\leq 0$, then there are significant effects of a monetary policy shock to the CPI and Industry, significant effects of an information shock to the REER, unemployement rate and the long term rate and significant effects of the spread shock to the REER, the unemployement rate and industrial production.
6 Conclusion
We investigated the presence of non-linearities in the impulse responses of EU countries that are not part of the EU exchange rate mechanism to exogenous monetary policy shocks. Using the strategy of Fanelli and Marsi [2022], we identify three shocks, a pure monetary shock, a central bank information shock and a spread shock. Focusing on their impact on the real exchange rate, the unemployment rate, the consumer price index, an index of industrial production and the long term government bond rate, we estimate benchmark linear impulse response functions and conditional and unconditional impulse response of non-linear functions of the shock using the method of Gonçalves et al. [2021]. The linear specification finds significant impacts of the pure monetary shock on all three nominal variables, while the spread shock impacts the two real variables. We find that, when compared with the sign effect specification, the linear specification underestimates the response of the long term rate to monetary policy, the unemployment rate to the information shock and the REER, CPI and industrial production to the spread shock. The size effect specification estimates impulse response that are close to the linear specification. We then perform a battery of significance test on both the coefficient of the non-linear transformation of the shock and both the conditional and unconditional impulse response functions. We find evidence suggesting the presence of sign non-linearities in the response to these monetary policy shocks. The evidence for size non-linearities is much weaker.
Appendix
Appendix A: AIC selection Results
In this section we provide the full model selection results with the information criterion of Akaike [1973]. The columns represent the horizons, while the first three rows are the optimal lag selection q, p, r. The last row is the optimal lag selection. That is, $t^{2}\equiv I_{1}=I_{2}=1$, $t\equiv I_{1}=1,I_{2}=0$ and $00\equiv I_{1}=I_{2}=$ 0.
Table 8: AIC Selection for Real Exchange Rate
h= 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
q 2 4 3 3 3 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 3
p 2 4 3 3 3 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 3
r 2 4 3 3 3 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
T 0 $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$
Table 9: AIC Selection for the Unemployment Rate
h= 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
q 2 2 2 2 2 2 2 2 3 3 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
p 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
r 2 2 2 2 2 2 2 2 3 3 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
T $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$
Table 10: AIC Selection for the Consumer Price Index
h= 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
q 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 3 2 2 2
p 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
r 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 3 2 2 2
T $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$
Table 11: AIC Selection for the Industrial production
h= 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
q 2 2 2 2 2 2 2 2 2 3 3 3 3 2 3 2 2 2 2 2 2 2 2 2 2
p 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
r 2 2 2 2 2 2 3 3 3 3 3 3 3 2 3 2 2 2 2 2 2 2 2 2 2
T $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t$ $t$ $t$ $t$ $t$
Table 12: AIC Selection for the Long Term Interest Rate
h= 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
q 2 2 2 7 7 7 7 7 7 7 7 7 7 7 7 7 7 7 7 6 5 4 4 4 3
p 2 2 2 2 6 6 5 5 3 2 2 2 2 2 2 2 2 2 2 2 2 2 4 4 2
r 2 2 2 3 7 7 7 7 4 2 2 7 7 7 5 4 3 3 2 3 3 2 2 2 2
T $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$
Appendix B: Comparaison with BIC
In this section we provide the full model selection results with the information criterion of Schwarz [1978]. The columns represent the horizons, while the first three rows are the optimal lag selection q, p, r. The last row is the optimal lag selection. That is, $t^{2}\equiv I_{1}=I_{2}=1$, $t\equiv I_{1}=1,I_{2}=0$ and $00\equiv I_{1}=I_{2}=$ 0.
Table 13: BIC Selection for the Real Exchange Rate
h= 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
q 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
p 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
r 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
T 0 0 0 $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$
Table 14: BIC Selection for the Unemployement Rate
h= 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
q 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
p 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
r 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
T $t$ $t$ $t^{2}$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t^{2}$
Table 15: BIC Selection for the Consumer Price Index
h= 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
q 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
p 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
r 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
T 0 $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$
Table 16: BIC Selection for the Industrial Production
h= 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
q 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
p 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
r 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
T $t$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$ $t$
Table 17: BIC Selection for the Long Term Rate
h= 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24
q 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
p 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
r 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2 2
T 0 $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$ $t^{2}$
Appendix C: Symmetry of the Shocks
In this section we plot the histogram of every shock and provide some summary statistics as well as the result of the symmetry tests of Cabilio and Masaro [1996] (CM), Miao et al. [2006] (M1) and Mira [1999] (M2). All these tests have an alternative hypothesis of asymmetry of the population distribution.
Figure 6: Histograms of the Three Shocks
Monetary Policy Shock Information Shock Spread Shock
Mean 0.0077 -0.0349 0.0284
Standard Deviation 0.9186 0.9751 0.8906
Skewness -0.2338 -0.3155 -0.7736
$80^{th}$ Quantile 0.539 0.447 0.395
CM Test Statistic 0.1640 -0.6955 0.6203
CM p-value 0.888 0.526 0.564
M1 Test Statistic 0.2045 -0.9095 0.9005
M1 p-value 0.896 0.478 0.554
M2 Test Statistic -0.1613 -0.6517 0.5599
M2 p-value 0.88 0.526 0.64
Table 18: Caption
The monetary policy and information shock are approximately symmetric as their skewness is small. The spread shock is a bit less symmetric due to the presence of many high magnitude negative shocks. However, all three tests fail to reject that the sample comes from a symmetric distribution. All three shocks were standardized and then assigned to a monthly time series therefore, their means are close to 0 and their standard deviations are close to 1.
Appendix D: IRFs and Inference for the pre-covid period
In this section we provide the estimated impulse response functions and the inference results for a sub-sample that ends in December 2019. That is:
Country Entry Date Exit Date
Czechia 2002M01 2019M12
Hungary 2002M01 2019M12
Poland 2002M01 2019M12
Romania 2005M04 2017M09
Sweden 2002M01 2017M05
Table 19: Composition of the Panel for the pre-pandemic period
We start with the linear specification and move to both the sign and size non-linear specifications.
The Linear Model
Figure 7: Results from the linear specification (6)
[…]
The horizontal axis represents the months since the shock (h). The rows are in order: the real exchange rate, the unemployment rate, the consumer price index, the index of industrial production and the long term interest rate. The columns are in order: the pure monetary policy shock, the information shock and the spread shock. The real exchange rate, consumer price index and industrial production index are in log points, while the unemployment rate and long term interest rate are in percentages. The dark blue lines represent the 1 standard deviation confidence interval while the light blue are $90\%$. The standard errors are computed using the method of Arellano [1987] with the “Jackknife” $HC_{3}$, HCCME..
Figure 7, provides the estimated impulse response functions of the linear specification. The pure monetary policy shock has significant, but temporary, effects on the real exchange rate and the CPI. The negative effects on industrial production are only significant at the $68\%$ significance level. The information shock has an quick but temporary impact on the real exchange rate and a delayed (and less significant) impact on CPI. Lastly, the spread shock has a significant negative impact on the unemployment rate and positive impact on industrial production. Its impact on CPI is positive at first (increase in CPI) but in the long run, this impact becomes negative.
The Non-Linear Model, Sign Non-Linearities
Figure 8: Impulse responses of the sign specification à la Gonçalves et al. [2021]
[…]
The red line is the original linear impulse response function, the blue line is the Gonçalves et al. [2021] non-linear IRF. The horizontal axis represents the months since the shock (h). The rows are in order: the real exchange rate, the unemployment rate, the consumer price index, the index of industrial production and the long term interest rate. The columns are in order: the pure monetary policy shock, the information shock and the spread shock. The real exchange rate, consumer price index and industrial production index are in log points, while the unemployment rate and long term interest rate are in percentages.
In figure 8, we plot in blue our estimate of the unconditional impulse response function to a 1 standard deviation disturbance in the sign-effect non-linear specification. In red is the original linear impulse response function to a 1 standard deviation shock. When comparing both we see that for most graphs the two IRFs nearly coincide, implying for them that on average both specification will estimate very similar IRFs. However, the monetary shock on the long term rate, is a lot stronger (temporarily) starting from horizon 5, when estimated with the sign-specification. For the information shock, the linear model estimates an effect on unemployment and industrial production that is close to zero, while the non-linear model estimates a much larger and positive IRF for unemployment and a negative IRF for industrial production. This implies that the effect of the information shock on real variables contains a sign effect. Lastly, the non-linear model estimates a larger, but temporary effect of the spread shock on industrial production, the real exchange rate and the consumer price index.
Figure 9: Conditional IRFs of the sign-effect specification
[…]
The rows are in order: the real exchange rate, the unemployment rate, the consumer price index, the index of industrial production and the long term interest rate. The columns are in order: the pure monetary policy shock, the information shock and the spread shock. The real exchange rate, consumer price index and industrial production index are in log points, while the unemployment rate and long term interest rate are in percentages.
Figure 9, compares two conditional impulse response functions estimated with the sign-effect specification. These are:
[…]
The blue IRFs are very similar to the unconditional Gonçalves et al. [2021] IRFs as the standard deviation and the estimated $\hat{A}$ are both close to 1. We can see that for monetary policy shocks, a (conditional) positive shock has a large negative effect on REER and a large positive effect on the long term rate, while a negative shock has a more restrained effect on both. For industrial production, it is the opposite, negative monetary policy shocks have more negative impacts. Positive information shocks have stronger permanent effects on unemployment, while negative shocks have stronger negative impacts on the long term rate. For the spread shock the main difference is in its effect on the unemployment rate. A negative spread shock, has an effect on the unemployment rate close to 0. On the other hand, positive shocks have large negative (decrease) on the unemployment rate. All other conditional Impulse response functions are similar when comparing negative and positive shocks.
The Non-Linear Model, Size Non-Linearities
Figure 10: Impulse responses of the size specification à la Gonçalves et al. [2021]
[…]
The rows are in order: the real exchange rate, the unemployment rate, the consumer price index, the index of industrial production and the long term interest rate. The columns are in order: the pure monetary policy shock, the information shock and the spread shock. The real exchange rate, consumer price index and industrial production index are in log points, while the unemployment rate and long term interest rate are in percentages.
Figure 10 plots the Gonçalves et al. [2021] unconditional impulse response functions for the size-effect specification in blue. We set the threshold $\bar{b}$ for each shock to be such that $P(|x_{t}|\leq\bar{b})\approx 0.6$. The shock used in computing of $\hat{A}$ is still 1 standard deviation. In red is the original linear impulse response function. The monetary policy shock has a slightly larger temporary effect on the REER for horizon 13 to 22. The information shock has a smaller effect on the unemployment rate, CPI and Industrial production when it is estimated with the non-linear specification. For all other impulse response functions, the linear and non-linear specification coincide. This suggests that there are no important size effects for a 1 standard deviation shock.
Figure 11: Impulse responses of the size specification à la Gonçalves et al. [2021]
[…]
The magenta, orange, red, green and blue line are for shocks that are 0.5, 0.75, 1, 1.25 and 1.5 times the standard deviation respectively. The rows are in order: the real exchange rate, the unemployment rate, the consumer price index, the index of industrial production and the long term interest rate. The columns are in order: the pure monetary policy shock, the information shock and the spread shock. The real exchange rate, consumer price index and industrial production index are in log points, while the unemployment rate and long term interest rate are in percentages.
Figure 11 presents an alternative way of illustrating the size effect using the method of Gonçalves et al. [2021]. Each line is an impulse response of the following form:
$IRF(h,a)=\frac{1}{a}\left(\hat{\psi}_{j,h,0}*a*\sigma+\sum_{t=1}^{T}\frac{1}{T}\left[f(x_{t}+a*\sigma)-f(x_{t})\right]*a*\hat{\Gamma}_{j,h,0}\right)$
Where $a\in\{0.5,0.75,1,1.25,1.5\}$. We can see that for many impulse responses the shape depends greatly on the size of the shock. This suggest that even if we did not see any differences in the impulse response between the linear and non-linear specification for the average shock (1 standard deviation), it could appear for different size shocks. For example, monetary policy shocks on the REER have bigger marginal effects for larger shocks for the first year ($h<12$), but smaller effects for the second year. The information shock’s marginal effect on REER and Unemployment is larger for smaller shocks. Lastly, the spread shock’s marginal effect on the REER, the CPI and Industrial Production is larger for smaller magnitude shocks. Therefore, if there are size effects, they are concave in the magnitude of the shock.
Inference Result
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 20: Significance of the Non-Linear Component of the Sign Effect Specification
[…]
We report the significance level of the estimated coefficient $\hat{\Gamma}_{j,h,0}$. The non-linear transformation is the absolute value of the shocks. White cells refer to p-values greater than 0.1, yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 20 reports the results of Wald tests (Wald [1943]) of the significance of the coefficient on the non-linear transformation of the shocks. As in Caravello and Martinez-Bruera [Working Paper, 2024], the results of these tests allow us to establish the presence of non-linearities at horizon h. We find multi-horizon significant sigh-effects of every type of shock on every outcome variable with the exception of the information shock on CPI. We describe the structure of the test bellow as we will present results on different version of it. We test the hypotheses,
$H_{0}^{1,h}:\hat{\Gamma}^{monetary}_{j,h,0}=0\quad H_{0}^{2,h}:\hat{\Gamma}^{information}_{j,h,0}=0\quad H_{0}^{3,h}:\hat{\Gamma}^{spread}_{j,h,0}=0$
To do so we conduct the Wald test:
$\hat{\beta}_{j,h}=$ $\langle\hat{\psi}^{monetary}_{j,h,0},\hat{\psi}^{information}_{j,h,0},\hat{\psi}^{spread}_{j,h,0},\hat{\Gamma}^{monetary}_{j,h,0},\hat{\Gamma}^{information}_{j,h,0},\hat{\Gamma}^{spread}_{j,h,0},...\rangle$
$R_{1}=$ $\langle 0,0,0,1,0,0,...\rangle,\quad R_{2}=\langle 0,0,0,0,1,0,...\rangle,\quad R_{3}=\langle 0,0,0,0,0,1,...\rangle$
$W_{i,j,h}=$ $(R_{i}\hat{\beta_{j,h}}^{\prime})^{\prime}(R_{i}\hat{\Omega}R_{i}^{\prime})^{-1}(R\hat{\beta_{j,h}}^{\prime})$
[…]
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 21: Significance of the Non-Linear Component of the Size Effect Specification
[…]
We report the significance level of the estimated coefficient $\hat{\Gamma}_{j,h,0}$. The non-linear transformation is $f(x_{t})=\mathbf{I}\{x_{t}\leq-\bar{b}\}(x_{t}+\bar{b})+\mathbf{I}\{x_{t}\geq\bar{b}\}(x_{t}-\bar{b})$ with $\bar{b}$ chosen such that $P(|x_{t}|<\bar{b})=0.6$. White cells refer to p-values greater than 0.1, yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 21 reports the result of the same Wald test as in table 2, but on the size effect specification. That is:
$R_{1}=$ $\langle 0,0,0,1,0,0,...\rangle,\quad R_{2}=\langle 0,0,0,0,1,0,...\rangle,\quad R_{3}=\langle 0,0,0,0,0,1,...\rangle$
The results are more muted then with the sign effect. For example, there does not seem to be any significant (and persistent over many horizon) size effect of any shock on the long term rate. The information shock has no size effect on industrial production. The persistent and significant non-linearities are that of the spread shock on CPI and Industry and all three shocks on the REER.
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24 25
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 22: Sign-Effect Impulse Response Function à la Gonçalves et al. [2021]
[…]
We report the significance level of the Gonçalves et al. [2021] IRF with the sign effect specification. White cells refer to p-values greater than 0.1, yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 22, reports our attempt at testing the significance of the Gonçalves et al. [2021] with a Wald test. This corresponds to the the following three hypotheses at each horizon,
$H_{0}^{i,h}:\hat{\psi}_{j,h,0}^{i}\delta^{i}+\hat{A}^{i}\hat{\Gamma}_{j,h,0}^{i}=0,\quad i\in\{monetary,information,spread\}$
To do so, we construct the following matrix of linear restrictions:
$R_{1}=$ $\langle\delta^{monetary},0,0,\hat{A}^{monetary}(\delta^{monetary}),0,0,...\rangle$
$R_{2}=$ $\langle 0,\delta^{information},0,0,\hat{A}^{information}(\delta^{information}),0,...\rangle$
$R_{3}=$ $\langle 0,0,\delta^{spread},0,0,\hat{A}^{spread}(\delta^{spread}),...\rangle$
Where: $\delta^{i}$ is the sample standard deviation of shock $i$ and $\hat{A}^{i}=\frac{1}{T}\sum_{t=1}^{T}\left[f(x_{t}^{i}+\delta^{i})-f(x_{t}^{i})\right]$, and recall that for the sign effect specification $f(x)=|x|$. The Wald Statistic is then constructed in the same way as before. This procedure is not exact as it does not take into account the estimation uncertainty in $\hat{A}^{i}$. Although we previously established the presence (or not) of non-linearities, the goal of this exercise is to see if these non-linearities do not “cancel out” the linear effect such that the weighted sum of both is not significant. The Gonçalves et al. [2021] IRF with the sign effect specification is significant for monetary shocks to the REER and the long term rate and the spread shock to CPI, unemployment and industrial production. The sign and magnitude of the impulse response is reported in figure 8.
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24 25
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 23: Size-Effect Impulse Response Function à la Gonçalves et al. [2021]
[…]
We report the significance level of the Gonçalves et al. [2021] IRF with the size effect specification. White cells refer to p-values greater than 0.1, yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 23 reports results of the same test as in table 4, albeit on the size-effect specification, with $\bar{b}$ such that $P(|x_{t}|<\bar{b})=0.6$. That is we test the hypothesis:
$H_{0}^{i,h}:\hat{\psi}_{j,h,0}^{i}\delta^{i}+\hat{A}^{i}\hat{\Gamma}_{j,h,0}^{i}=0,\quad i\in\{monetary,information,spread\}$
The wald statistic is constructed in the same way as in table 4. As when we tested the significance of $\hat{\Gamma}_{j,h,0}^{i}$ alone, we find more muted results. The impulse response function is only significant (for multiple horizons) for monetary shocks to REER and CPI.
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 24: Conditional IRFs Positive Sign-Effect
[…]
We report the significance level of the conditional IRFs for a positive shock with the sign effect specification. White cells refer to p-values greater than 0.1, yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 24, provides the results for significance tests on the sign-effect, positive-shock, conditional impulse response functions. That is we test the hypotheses:
$H_{0}^{i,h}:\hat{\psi}_{j,h,0}^{i}+\hat{\Gamma}_{j,h,0}^{i}=0,\quad i\in\{monetary,information,spread\}$
To do so we construct the restriction matrix:
$R_{1}=\langle 1,0,0,1,0,0,...\rangle\quad R_{2}=\langle 0,1,0,0,1,0,...\rangle\quad R_{3}=\langle 0,0,1,0,0,1,...\rangle$
We then construct the Wald statistic in the same way. We find that conditional on $x_{t}\geq 0$ and $\delta>0$, there are significant effects of the monetary policy shock to REER, CPI and the Long term Rate, significant effects of the information shock to unemployment at horizon 11 to 18 and significant effects of the spread shocks to CPI, Industry and Unemployment.
h= 0 1 2 3 4 5 6 7 8 9 10 11 12
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
h= 13 14 15 16 17 18 19 20 21 22 23 24
REER Monetary
Information
Spread
Unemp. Monetary
Information
Spread
CPI Monetary
Information
Spread
Industry Monetary
Information
Spread
LT Rate Monetary
Information
Spread
Table 25: Conditional IRFs Negative Sign-Effect
[…]
We report the significance level of the conditional IRFs for a negative shock with the sign effect specification. Yellow cells are p-values between 0.1 and 0.05 and green cells are p-values smaller than 0.05.
Table 25, provides the results for significance tests on the sign-effect, negative-shock, conditional impulse response functions. That is we test the hypotheses:
$H_{0}^{i,h}:\hat{\Gamma}_{j,h,0}^{i}-\hat{\psi}_{j,h,0}^{i}=0,\quad i\in\{monetary,information,spread\}$
Which imply the restriction matrices:
$R_{1}=\langle-1,0,0,1,0,0,...\rangle\quad R_{2}=\langle 0,-1,0,0,1,0,...\rangle\quad R_{3}=\langle 0,0,-1,0,0,1,...\rangle$
From the wald test we find that if $\delta<0$ and $x_{t}\leq 0$, then there are significant effects of a monetary policy shock to the CPI and Industry, significant effects of an information shock to the REER, unemployement rate and the long term rate and significant effects of the spread shock to the REER, the CPI and industrial production.